I need to find the equation for the line, in general from, with the given properties.
(-8,1) (-6,-2)
i do have the answer, but i don't know how to solve it. I guessed the answer.

Answer by stanbon(75887)

find the equation for the line, in general from, with the given properties.
(-8,1) (-6,-2)
-------
slope = (1--2)/(-8--6)
= 3/-2
= -3/2
------
Form: y = mx + b
solve for "b":
1 = (-3/2)(-8) + b
1 = 12 + b
b = -11
----
Equation:
y = (-3/2)x - 11
========================
cheers,
Stan H.
========================